OpenStudy (anonymous):

I missed 4 questions on this test. There are 20 questions total. What is the ratio of incorrect answers to correct answers? The correct answer is expressed in its simplest form. 4/20 1/5 4/16 1/4

OpenStudy (anonymous):

1/5

OpenStudy (bibby):

First set up the proportion.\[\frac{ incorrect }{ correct } = \frac{ 4 }{ 20-4 }\]

OpenStudy (anonymous):

wich equals 1/5 right

OpenStudy (anonymous):

Can u help me next bibby with my new open question

OpenStudy (bibby):

This isn't asking incorrect/total questions. Always reread the question twice. 4/20 = 1/5. but you didn't get 20 correct. you got 16 correct. as such your ratio is 4/16

OpenStudy (bibby):

uh yeah. message me next time and don't hijack other peoples questions

OpenStudy (anonymous):

Sorry im new to this site

OpenStudy (anonymous):

wait so 1/5 isn't the answer it 4/16 ?

OpenStudy (bibby):

Yeah. It's a misleading question. \[\frac{ incorrect answers }{ correct answers } = \frac{ 4 }{ 20-4 } = \frac{ 4 }{ 16 } = \frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

o ok sorry about that

OpenStudy (bibby):

"everybody makes mistakes" - hannah montana

OpenStudy (anonymous):

lol ya